OpenStudy (anonymous):

Madison deposited $18 in a savings account that earns 2.5% simple interest. Which graph represent this scenario

OpenStudy (jdoe0001):

the graph that says Madison? =)

OpenStudy (anonymous):

Ok i don't how to upload the graph but can I tell you the points? @jdoe0001

OpenStudy (jdoe0001):

see the blue button that says [Attach File], next to the [Draw] button?

OpenStudy (anonymous):

yes i do

OpenStudy (anonymous):

Can you please give me one minute until i upload the graph @jdoe0001

OpenStudy (jdoe0001):

well... we know her original amount is $18 => principal we know the interest rate is 2.5% => \(\bf \frac{2.5}{100}\implies 0.025\) so.... to find how much she did after "t" years, just get the values say for example after \(\bf A=P+Prt\implies A=P(1+rt)\qquad t=years\\ \quad \\ \begin{array}{ccllll} years&Amount\\ \hline\\ {\color{red}{ 1}}&A=18[1+0.025\cdot {\color{red}{ 1}})]\\ {\color{red}{ 2}}&A=18[1+0.025\cdot {\color{red}{ 2}})]\\ {\color{red}{ 3}}&A=18[1+0.025\cdot {\color{red}{ 3}})]\\ {\color{red}{ 4}}&A=18[1+0.025\cdot {\color{red}{ 4}})]\\ {\color{red}{ 5}}&A=18[1+0.025\cdot {\color{red}{ 5}})]\\ ... \end{array}\)

OpenStudy (anonymous):

I'm looking for the interest rate so I use the same formula?

OpenStudy (jdoe0001):

well, you want to know how much is her total for each year, or A= Amount of balance so that'd be \(\bf A=P(1+rt)\)

OpenStudy (anonymous):

So the first year is 18.45?

OpenStudy (jdoe0001):

yeap

OpenStudy (jdoe0001):

2 => 18.9 3 => 19.35 and so on

OpenStudy (anonymous):

Ok so far I get it perfectly but I don't how see with the graphs?

OpenStudy (anonymous):

I'm sorry with graphs

OpenStudy (triciaal):

what are you calculating? Interest I varies directly with the time Interest I= PRT when t = o P. principal is 18

OpenStudy (jdoe0001):

try say 2, 6 and 10 for the years

OpenStudy (jdoe0001):

see if you can match up with one of the choices

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Yeah I think the one that makes more sense is graph 3 @jdoe0001

OpenStudy (jdoe0001):

yeap, the last graph you posted

OpenStudy (triciaal):



OpenStudy (anonymous):

Yes thank you so much @jdoe0001